Assume that A and B are matrices of the same size. Determine an expression for A if 2A − B = 5(A + 2B)


1
Expert's answer
2021-05-11T09:06:03-0400

Given expression is-


"2A-B=5(A+2B)

\\[9pt]

2A-B=5A+10B

\\[9pt]

3A=-11B

\\[9pt]

A=\dfrac{-11}{3}B"